# Polarization Purity and Dispersion Characteristics of Nested Antiresonant Nodeless Hollow-Core Optical Fiber at Near- and Mid-IR Wavelengths for Quantum Communications

### v1

@@ 5 Conclusion
Analyzing the polarization characteristics, we find that the polarization extinction ratio (PER) significantly influences the quantum bit error rate (QBER) and secure key rates, which are a crucial parameters for quantum communication protocols. For the BB84 protocol, it is apparent that the quality of the polarization-based quantum key distribution (QKD) is highly sensitive to the spectral bandwidth of the photons. With a QBER acceptable for secure quantum communication only at bandwidths approximately below 10 nm, this places a stringent limitation on the operability of a polarization-based BB84 protocol using this fiber, determined primarily by the polarization purity that the fiber can maintain. On the other hand, the BBM92 protocol, which utilizes entangled photons and allows for synchronization in a continuous-wave mode, seems more promising for implementation of polarization-based QKD with this fiber. The PER measurements and the subsequent analysis indicate that the fiber could sustain a robustly positive asymptotic secret key rate, thereby reinforcing the potential of the 2-$\mu$m window for both prepare-and-measure and entanglement-based quantum communication protocols. We note that fibers are typically not considered ideal for polarization-based QKD due to inherent challenges, giving preference for using the temporal degree of freedom, such as time-bin encoding which benefits from low-latency propagation in air. However, the advent of NANF fibers also opens the possibility of not excluding polarization as a viable degree of freedom for long distance quantum communication.
Our findings highlight the critical parameters that must be optimized for the development of reliable and secure quantum communication systems. Furthermore, the results point to further research into the deployment of such fibers in a broader range of quantum technologies, possibly opening new avenues for quantum communication at higher data rates, over longer distances, and under more practical conditions.
@@
Funding D.F. is supported by the Royal Academy of Engineering under the Chairs in Emerging Technologies scheme. M.C. acknowledges support from the UK Research and Innovation (UKRI) and the UK Engineering and Physical Sciences Research Council (EPSRC) Fellowship (“In-Tempo” EP/S001573/1). A.C.D. acknowledges support from the EPSRC, Impact Acceleration Account (EP/R511705/1).

### v2

@@ 5 Conclusion
Analyzing the polarization characteristics, we find that the polarization extinction ratio (PER) significantly influences the quantum bit error rate (QBER) and secure key rates, which are a crucial parameters for quantum-communication protocols. For the BB84 protocol, it is apparent that the quality of the polarization-based quantum key distribution (QKD) is highly sensitive to the spectral bandwidth of the photons. With a QBER acceptable for secure quantum communication only at bandwidths approximately below 10 nm, this places a stringent limitation on the operability of a polarization-based BB84 protocol using this fiber, determined primarily by the polarization purity that the fiber can maintain. On the other hand, the BBM92 protocol, which utilizes entangled photons and allows for synchronization in a continuous-wave mode, seems more promising for implementation of polarization-based QKD with this fiber. The PER measurements and the subsequent analysis indicate that the fiber could sustain a robustly positive asymptotic secret key rate, thereby reinforcing the potential of the 2-$\mu$m window for both prepare-and-measure and entanglement-based quantum-communication protocols. We note that fibers are typically not considered ideal for polarization-based QKD due to inherent challenges, giving preference for using the temporal degree of freedom, such as time-bin encoding which benefits from low-latency propagation in air. However, the advent of NANF fibers also opens the possibility of not excluding polarization as a viable degree of freedom for long distance quantum communication. These results in the 2-$\mu$m window also further open up the bandwith capabilities of HCFs. Indeed, recent research [41] has demonstrated the capability of NANFs to transmit single-photon and classical optical channels separated by hundreds of nanometers with a remarkable 30 dB reduction in background noise compared to using closer wavelengths. The use of HCFs in this context represents a promising advancement in the development of practical and scalable quantum networks that can coexist with classical communication infrastructure.
Our findings highlight the critical parameters that must be optimized for the development of reliable and secure quantum-communication systems. Furthermore, the results point to further research into the deployment of such fibers in a broader range of quantum technologies, possibly opening new avenues for quantum communication at higher data rates, over longer distances, and under more practical conditions.
@@
Funding D.F. is supported by the Royal Academy of Engineering under the Chairs in Emerging Technologies scheme and the Engineering and UK Physical Sciences Research Council (EPSRC) project Quantic (EP/T00097X/1). M.C. acknowledges support from the UK Research and Innovation (UKRI) and the EPSRC Fellowship (“In-Tempo” EP/S001573/1). A.C.D. acknowledges support from the EPSRC Impact Acceleration Account (EP/R511705/1).
@@
[…]
Disclosures The Authors declare no conflicts of interest.
[…]
Data Availability All the data supporting the conclusions reported in this manuscript are available online [http://dx.doi.org/10.5525/gla.researchdata.1772]. Additional data related to this paper may be requested from the authors.